# Noisy Quantum Learning Theory

### v1

@@ Appendix F Pauli Tomography in Noisy Quantum Learning Theory
In this section, we construct a quantum state-discrimination task that is strictly easier than performing Pauli shadow tomography [CGY24]. We show that any quantum algorithm restricted to single-copy measurements requires at least $(2/f(\lambda))^{n}$ measurements to succeed with high probability, where $f(\lambda)\in[0,1]$ depends inversely on the noise rate. More generally, even with $k$ qubits of quantum memory, any learning strategy still needs $2^{\,n-k}(1-\lambda)^{-n}$ samples. These bounds immediately yield lower bounds for noisy Pauli shadow tomography. Since learning Pauli observables (or stabilizer states [Mon17]) is the canonical application of Bell sampling, our information-theoretic lower bounds — which require only a single invocation of the noise channel — demonstrate that Bell-measurement-based strategies are exponentially degraded by local noise.
We then give a single-copy algorithm for noisy Pauli shadow tomography and a two-copy algorithm which can solve the state-discrimination task. The latter has sample complexity matching the lower bound for $k=n$ qubits of quantum memory up to nonleading factors of $n$. This establishes a polynomial separation, dependent on the noise rate, between our two-copy algorithm and any single-copy protocol. As the noise rate tends to zero, this polynomial advantage becomes exponential, recovering the ideal quantum advantage for Pauli shadow tomography. We begin by defining the broader learning problem.
@@
This physically motivated learning task has been extensively studied, and optimal bounds on its sample complexity in the noiseless setting are given in [CGY24]. Here, we treat this problem in the context of NBQP-type errors. Our bounds neglect the well-established $\varepsilon$-dependence of this task for clarity and to emphasize the interplay between noise rate and instance size, as $\varepsilon$-dependence is not altered by noise. For our lower bounds, we consider three models of $\lambda$-noisy quantum algorithms:
@@ F.1 Lower bound for single-copy measurements without quantum memory
Before proceeding, we state two lemmas characterizing the operator norm of SWAP operators under the depolarizing channel, whose proofs we defer to Appendix G.2.
@@
Both lower bounds must hold for all $\lambda\in[0,1]$, so our final lower bound is the maximum of the two. Note that for $\lambda\rightarrow 1$, our sample complexity bound now goes to infinity, while for $\lambda\rightarrow 0$, our bound goes to $2^{n}$, matching the noiseless analysis in [CCHL22]. ∎
@@ F.2 Lower bound with $k$ qubits of memory and constant queries per experiment
The next model we consider is a $\lambda$-noisy algorithm with access to $k<n$ qubits of quantum memory, but which can only make a fixed number of queries $c$ to the oracle in each experiment before discarding and re-initializing its quantum register. This is analogous to a noisy quantum register which can only interact with the unknown state a constant number of times before projecting and recording a classical outcome, so we say the memory qubits have a limited lifetime $c$. This framework was introduced in [CGY24], and has a learning tree representation characterized as follows.
@@ F.3 Lower bound with $k$ qubits of memory and unbounded depth
For $T<\Omega(2^{(n-k)/3}(1-\lambda)^{n/3})$, we find that the total variation distance is $o(1)$. Hence, to achieve a success probability $\geq 2/3$, Lemma C.11 requires that $T\geq\Omega(2^{(n-k)/3}(1-\lambda)^{n/3})$. As $\lambda\rightarrow 1$, our lower bound becomes infinite, which is expected because the I+P state becomes maximally mixed. As $\lambda\rightarrow 0$, we recover the $\Omega(2^{(n-k)/3})$ lower bound from [CCHL22],indicating that our bounds are tight up to constants in the exponents. ∎
Previous memory-aware lower bounds in [CCHL22] became trivial once $k=n$. Note here that even when $k=n$, the sample complexity lower bound still scales as $\Omega((1-\lambda)^{-n/3})$; hence, no sample efficient two-copy algorithm (including any strategy leveraging Bell measurements) for Pauli shadow tomography, or even the easier many-vs.-one discrimination problem, can exist.
@@ F.4 Single-copy noisy strategy
Here we provide an algorithm using only single copies to solve Pauli shadow tomography in the presence of NBQP noise. To align with our lower bounds, we consider the setting in which $\varepsilon=\delta=1/3$; the dependence on these parameters in the sample complexity are the standard $\varepsilon^{-2},\log(1/\delta)$. A simple approach is to apply classical shadow tomography [HKP20]. We first recall the algorithm.
@@ 8   Measure $U^{\dagger}\rho U$ in the computational basis to obtain $n$-bistring $s$
The subroutine $\textsc{MedianOfMeans}(S,O,K)$ simply batches the set $S$ into $S/K$ nonoverlapping sets of size $K$, computes $\mathsf{tr}(O\mathcal{M}^{-1}(s))$ for each $\mathcal{M}^{-1}(s)$ stored in each batch, and returns the median of the means of all batches for the estimator $\hat{O}_{i}$. It is also generally required that $\mathcal{E}$ matches moments of the Haar measure. For details, see [HKP20].
@@
To simplify notation, we define the noisy scrambling channel $\mathcal{C}_{U,\lambda}[\cdot]=\mathcal{D}_{\lambda}^{\otimes 2n}[U\mathcal{D}_{\lambda}^{\otimes 2n}[\cdot]U^{\dagger}]$. Here, $U$ will be a product of Haar-random single qubit unitaries, and the output of the channel represents the quantum state directly before computational-basis measurement in the classical shadows algorithm. The ideal, noiseless reconstruction channel corresponding to a fixed $U$ is given by $\mathcal{R}_{U}[\cdot]=U\cdot U^{\dagger}$. Recall that in classical shadow tomography [CW20, HKP20], reconstruction of expectation values from measurements is performed classically, and so we do not incur errors in this step.
With this notation in hand, we state known results on the shadow norm of Pauli observables.  [HCY23] and  [BKGJ24] demonstrate that when the ensemble $\mathcal{E}$ of quantum channels applied to the given state is invariant under local rotations (which holds trivially for the single-qubit Haar random ensemble and depolarizing channel), the classical shadows channel $\mathcal{M}[\rho]$ as defined in Algorithm 1 has the following diagonal representation in the Pauli basis:
@@
We remark that this result can be generalized slightly in the case where the classical shadows algorithm is implemented by a 1-dimensional brickwork circuit of depth $D$. Ref. [HGM+25] shows that for such circuits (layered with interstitial depolarizing noise of strength $\lambda$),
@@ F.5 Two-copy noisy Bell sampling and quantum-enhanced advantage
Note that our lower bound for the model of $k$ qubits of quantum memory has a noise dependence that scales as $(1-\lambda)^{-n}$, up to a constant in the exponent. Hence, all two-copy strategies for Pauli shadow tomography are irrecoverably degraded by just the single layer of noise considered in our lower bounds, including all strategies which utilize Bell measurement. In this section, we give a Bell-measurement based algorithm for Dec-IP which, up to constants in the exponent and a nonleading polynomial factor in $n$, matches this exponential scaling in $\lambda$.
@@ 13   With $q\in\mathbb{F}_{2}^{2n}$ denoting the symplectic bitstring of $Q$, define
Classically enumerating over all non-identity Paulis,w then define our estimator $\hat{Z}_{Q}$ to be the empirical mean of $X_{Q}$. In expectation, only $\mathbb{E}[\hat{Z}_{P}]$ is greater than 0, so our algorithm chooses the largest empirical $\hat{Z}_{Q}$ as the final estimator. We note that, under success, our algorithm has also identified $P$ correctly, and is thus stronger than simply solving Dec-IP. It is plausible that with an appropriate reframing, e.g. fixing a Pauli $P$ in the alternate hypothesis as in Ref. [HFP22], we may obtain a tighter gap between our lower and upper bounds.
@@
Appendix G Deferred Proofs
G.1 Proof of Lemma D.9

### v2

@@ Appendix F Pauli Tomography in Noisy Quantum Learning Theory
In this section, we construct a quantum state-discrimination task that is strictly easier than performing Pauli shadow tomography [weiyuan_paulis]. We show that any quantum algorithm restricted to single-copy measurements requires at least $(2/f(\lambda))^{n}$ measurements to succeed with high probability, where $f(\lambda)\in[0,1]$ depends inversely on the noise rate. More generally, even with $k$ qubits of quantum memory, any learning strategy still needs $2^{\,n-k}(1-\lambda)^{-n}$ samples. These bounds immediately yield lower bounds for noisy Pauli shadow tomography. Since learning Pauli observables (or stabilizer states [montanaro2017learningstabilizerstatesbell]) is the canonical application of Bell sampling, our information-theoretic lower bounds — which require only a single invocation of the noise channel — demonstrate that Bell-measurement-based strategies are exponentially degraded by local noise.
We then give a single-copy algorithm for noisy Pauli shadow tomography and a two-copy algorithm which can solve the state-discrimination task. The latter has sample complexity matching the lower bound for $k=n$ qubits of quantum memory up to nonleading factors of $n$ and constants in the exponents. This establishes a polynomial separation, dependent on the noise rate, between our two-copy algorithm and any single-copy protocol. As the noise rate tends to zero, this polynomial advantage becomes exponential, recovering the ideal quantum advantage for Pauli shadow tomography. We begin by defining the broader learning problem.
@@
This physically motivated learning task has been extensively studied, and optimal bounds on its sample complexity in the noiseless setting are given in [weiyuan_paulis]. Here, we treat this problem in the context of NBQP-type errors. Our bounds neglect the well-established $\varepsilon$-dependence of this task for clarity and to emphasize the interplay between noise rate and instance size, as $\varepsilon$-dependence is not altered by noise. For our lower bounds, we consider three models of $\lambda$-noisy quantum algorithms:
@@ F.1 Lower bound for single-copy measurements without quantum memory
Before proceeding, we state two lemmas characterizing the operator norm of SWAP operators under the depolarizing channel, whose proofs we defer to Appendix H.2.
@@
Both lower bounds must hold for all $\lambda\in[0,1]$, so our final lower bound is the maximum of the two. Note that for $\lambda\rightarrow 1$, our sample complexity bound now goes to infinity, while for $\lambda\rightarrow 0$, our bound goes to $2^{n}$, matching the noiseless analysis in [chen2021exponentialseparationslearningquantum]. ∎
@@ F.2 Lower bound with $k$ qubits of memory and constant queries per experiment
The next model we consider is a $\lambda$-noisy algorithm with access to $k<n$ qubits of quantum memory, but which can only make a fixed number of queries $c$ to the oracle in each experiment before discarding and re-initializing its quantum register. This is analogous to a noisy quantum register which can only interact with the unknown state a constant number of times before projecting and recording a classical outcome, so we say the memory qubits have a limited lifetime $c$. This framework was introduced in [weiyuan_paulis], and has a learning tree representation characterized as follows.
@@ F.3 Lower bound with $k$ qubits of memory and unbounded depth
For $T<\Omega(2^{(n-k)/3}(1-\lambda)^{n/3})$, we find that the total variation distance is $o(1)$. Hence, to achieve a success probability $\geq 2/3$, Lemma C.11 requires that $T\geq\Omega(2^{(n-k)/3}(1-\lambda)^{n/3})$. As $\lambda\rightarrow 1$, our lower bound becomes infinite, which is expected because the I+P state becomes maximally mixed. As $\lambda\rightarrow 0$, we recover the $\Omega(2^{(n-k)/3})$ lower bound from [chen2021exponentialseparationslearningquantum],indicating that our bounds are tight up to constants in the exponents. ∎
Previous memory-aware lower bounds in [chen2021exponentialseparationslearningquantum] became trivial once $k=n$. Note here that even when $k=n$, the sample complexity lower bound still scales as $\Omega((1-\lambda)^{-n/3})$; hence, no sample efficient two-copy algorithm (including any strategy leveraging Bell measurements) for Pauli shadow tomography, or even the easier many-vs.-one discrimination problem, can exist.
@@ F.4 Single-copy noisy strategy
Here we provide an algorithm using only single copies to solve Pauli shadow tomography in the presence of NBQP noise. To align with our lower bounds, we consider the setting in which $\varepsilon=\delta=1/3$; the dependence on these parameters in the sample complexity are the standard $\varepsilon^{-2},\log(1/\delta)$. A simple approach is to apply classical shadow tomography [classical_shadow]. We first recall the algorithm.
@@ 8   Measure $U^{\dagger}\rho U$ in the computational basis to obtain $n$-bistring $s$
The subroutine $\textsc{MedianOfMeans}(S,O,K)$ simply batches the set $S$ into $S/K$ nonoverlapping sets of size $K$, computes $\mathsf{tr}(O\mathcal{M}^{-1}(s))$ for each $\mathcal{M}^{-1}(s)$ stored in each batch, and returns the median of the means of all batches for the estimator $\hat{O}_{i}$. It is also generally required that $\mathcal{E}$ matches moments of the Haar measure. For details, see [classical_shadow].
@@
To simplify notation, we define the noisy scrambling channel $\mathcal{C}_{U,\lambda}[\cdot]=\mathcal{D}_{\lambda}^{\otimes 2n}[U\mathcal{D}_{\lambda}^{\otimes 2n}[\cdot]U^{\dagger}]$. Here, $U$ will be a product of Haar-random single qubit unitaries, and the output of the channel represents the quantum state directly before computational-basis measurement in the classical shadows algorithm. The ideal, noiseless reconstruction channel corresponding to a fixed $U$ is given by $\mathcal{R}_{U}[\cdot]=U\cdot U^{\dagger}$. Recall that in classical shadow tomography [cotler2020quantum, classical_shadow], reconstruction of expectation values from measurements is performed classically, and so we do not incur errors in this step.
With this notation in hand, we state known results on the shadow norm of Pauli observables.  [hu2023classical] and  [bu2024classical] demonstrate that when the ensemble $\mathcal{E}$ of quantum channels applied to the given state is invariant under local rotations (which holds trivially for the single-qubit Haar random ensemble and depolarizing channel), the classical shadows channel $\mathcal{M}[\rho]$ as defined in Algorithm 1 has the following diagonal representation in the Pauli basis:
@@
We remark that this result can be generalized slightly in the case where the classical shadows algorithm is implemented by a 1-dimensional brickwork circuit of depth $D$. Ref. [hu2025demonstration] shows that for such circuits (layered with interstitial depolarizing noise of strength $\lambda$),
@@ F.5 Two-copy noisy Bell sampling and quantum-enhanced advantage
Note that our lower bound for the model of $n$ qubits of quantum memory has a noise dependence that scales as $(1-\lambda)^{-n}$, up to a constant in the exponent. Hence, all two-copy strategies for Pauli shadow tomography are irrecoverably degraded by just the single layer of noise considered in our lower bounds, including all strategies which utilize Bell measurement. In this section, we give a Bell-measurement based algorithm for Dec-IP which, up to constants in the exponent and a nonleading polynomial factor in $n$, matches this exponential scaling in $\lambda$.
@@ 13   With $q\in\mathbb{F}_{2}^{2n}$ denoting the symplectic bitstring of $Q$, define
Classically enumerating over all non-identity Paulis,w then define our estimator $\hat{Z}_{Q}$ to be the empirical mean of $X_{Q}$. In expectation, only $\mathbb{E}[\hat{Z}_{P}]$ is greater than 0, so our algorithm chooses the largest empirical $\hat{Z}_{Q}$ as the final estimator. We note that, under success, our algorithm has also identified $P$ correctly, and is thus stronger than simply solving Dec-IP. It is plausible that with an appropriate reframing, e.g. fixing a Pauli $P$ in the alternate hypothesis as in Ref. [huang2022foundations], we may obtain a tighter gap between our lower and upper bounds.
@@
Although local noise exponentially degrades Bell sampling, polynomial quantum-classical separations can persist and still yield substantial experimental gains [Chen_2024_pauli, seif2024, Oh_2024, Liu_2025]. For near-term experiments at fixed instance size $n$, the more relevant question is how small the noise rate $\lambda$ must be to achieve a desired advantage. Our Dec-IP analysis assumes $\lambda=\Theta(1)$, but if $\lambda=\Theta(1/n)$, the two-copy Bell-sampling strategy regains an $\exp(\Omega(n))$ separation over any single-copy protocol. Allowing $\lambda$ to scale with $n$ is not an asymptotic advantage statement, but it yields a concrete finite-size criterion: when $\lambda\ll 1/n$, the sample-complexity gap can grow essentially exponentially with $n$. Bounds like Corollary 2.7 make this dependence explicit, translating target performance into quantitative noise requirements.
Appendix G A Threshold for Heisenberg-Limited Metrology
In this Section, we demonstrate that the presence of interstitial noise in quantum-enhanced experiments has fundamental implications for quantum metrology. Here, we work with standard definitions for single-parameter quantum sensing.
Definition G.1 (Single-parameter, single-probe quantum sensing task).
Consider an $n$-qubit Hamiltonian $H=\omega G$, where $\omega$ is a real-valued scalar, and a set of Lindblad jump operators $\{L_{k}\}$ of size $m$. Moreover, fix a precision $\varepsilon>0$. Given a single quantum probe $\rho_{P}$ on $n$ qubits, evolving according to the time-homogeneous Lindblad master equation
$\frac{d\rho}{dt}=-i[H,\rho]+\sum_{j=1}^{m}L_{j}\rho L_{j}^{\dagger}-\frac{1}{2}\{L_{j}^{\dagger}L_{j},\rho\}\ ,$ (173)
the task of single-parameter quantum sensing is to estimate $\omega$ to within absolute precision $\varepsilon$. Ancillary quantum memory and quantum information processing may be used.
Strategies that do not require coherent quantum information processing can perform single-parameter sensing using a total sensor evolution time $t=O(\varepsilon^{-2})$, known as the Standard Quantum Limit (SQL). However, the gold standard is to achieve the so-called Heisenberg limit (HL) of $t=O(\varepsilon^{-1})$. We note that an HL sensing protocol using total time $t$ and a single probe can often be reinterpreted as an $N$-probe protocol requiring sensing time $t/N$ per probe, whereas for SQL protocols, using $N$ parallel sensors will require $t/\sqrt{N}$ sensing time.
A well-known result in metrology, due to [Zhou_2018], proposes a necessary and sufficient condition for the achievability of HL metrology. Namely, they define the Hamiltonian-Not-in-Lindblad-Span (HNLS) criterion, which requires that $H$ not be contained in the linear span of $I,L_{j},L_{j}^{\dagger},L_{j}^{\dagger}L_{k}$ for all Lindblad jumps and their pairwise products. If HNLS holds, Ref. [Zhou_2018] demonstrates the existence of a quantum error-correcting code which can correct for jumps while preserving logical time evolution under the signal Hamiltonian (when executed using fast quantum control and between short-time queries to the signal). Conversely, they prove that when HNLS is violated, there is no protocol which can outperform the SQL.
At a high level, the error-correction scheme in [Zhou_2018] is enabled by the assumption that quantum error-correction is effectively instantaneous, so that the dissipator, which only evolves for short times, is unlikely to generate high-weight errors consisting of multiple jump terms. In this regime, error-correction handles first-order-in-time contributions from the dissipator, while higher weight errors occur with exponentially low probability. In light of our results, we can make several practical refinements. First, real-world hardware can experience noise at the interface between sensor and computing degrees of freedom, whenever the protocol alternates between signal queries and error-correction rounds. Moreover, quantum control takes finite time, and any noise that is native to the quantum computer, which may not be captured in the signal Lindbladian, can still accrue. The strength of this interface-level noise channel is thus parametrized both by the native noise rate of the quantum computer and the time each round of quantum control takes. As a result, the continuous-time approximations made in [Zhou_2018], where noise and error correction occur as simultaneous processes, require refinement in a more practical setting.
To make this precise in an instructive but simple way, we assume each round of quantum control occupies a time at least $\tau$. Moreover, in line with NBQP, we assume that the “interface” noise channel which appears between signal queries and quantum control is depolarizing. To align with the Lindbladian formalism, we assume this noise is generated by the channel $e^{\mathcal{L}_{\textnormal{dep}}^{\eta}\tau}$, where
$\mathcal{L}_{\textnormal{dep}}^{\eta}(\rho)=\eta\sum_{j=1}^{n}\sum_{\sigma\in\{X,Y,Z\}}(\sigma^{(j)}\rho\sigma^{(j)}-\rho)$ (174)
is the Lindbladian which generates a local depolarizing channel. Here, we parametrize the depolarization strength by $\eta$ and $\tau$, noting that as quantum channels, $\exp(\mathcal{L}_{\textnormal{dep}}^{\eta}\tau)$ and $\mathcal{D}^{\otimes n}_{\lambda}$ are equivalent when $\lambda=1-\exp(-4\eta\tau)$. We then work with an effective description of interstitial noise, in which every round of quantum control is assumed to be a noiseless black box interleaved with $\exp(\mathcal{L}_{\textnormal{dep}}^{\eta}\tau)$.
If at every round of quantum control in the procedure from [Zhou_2018], we introduce this interface noise and assume HNLS is satisfied, how does the protocol’s HL sensitivity change?
Theorem G.2.
Consider a single-parameter sensing task satisfying the HNLS criterion, with interstitial depolarizing noise applied after each round of quantum control. There exist thresholds $\tilde{\varepsilon}(\eta,\tau,n)$ and $\tilde{t}(\eta,\tau,n)$ such that the protocol of Ref. [Zhou_2018] achieves Heisenberg-limited sensitivity for target precision $\varepsilon>\tilde{\varepsilon}$ and total evolution time $t<\tilde{t}$, and crosses over to the Standard Quantum Limit otherwise. For $n\eta\tau\ll 1$, the thresholds scale as $\tilde{t}=O(1/\eta n))$ and $\tilde{\varepsilon}=O(\eta n)$.
Note that because our bounds treat the quantum control as a black box, they extend to any protocol achieving Heisenberg-limited sensing in the absence of interface noise if the protocol uses many interleaved rounds of signal queries and quantum control, only making measurements at the end. This result places a fundamental ceiling on the achievability of the Heisenberg limit in the presence of experimental noise, absent other problem-dependent error-mitigation strategies. Without leveraging such strategies to correct for both signal noise and interface noise, HL sensitivity is preserved up to only inverse-polynomial total time in the number of qubits.
Our proof will elucidate the hardware dependence of HL sensitivity to the physical noise rate and the speed of error correction. This perspective is crucial to quantify precision limits and requisite experimental parameters for practical error-correction-based metrology protocols.
To proceed with the proof, we require some notation. In the language of [Zhou_2018], each round of sensing and error-correction (without interface-level noise) is described by the quantum channel $\Psi(\rho)=\mathcal{R}\circ e^{\mathcal{L}t_{single}}$, where $t_{1}$ is the time for a single round and $\mathcal{R}$ is a recovery channel which projects the joint state of probe and ancillas onto a chosen codespace which protects the Hamiltonian against Lindblad jumps. This high-level channel description will suffice for our argument, but the full construction of $\mathcal{R}$ is detailed in [Zhou_2018]. Upon incorporating NBQP-style noise, the channel becomes $\Phi(\rho)=\mathcal{R}\circ e^{\mathcal{L}_{dep}^{\eta}\tau}\circ e^{\mathcal{L}t_{1}}$. The entire protocol consists of $N$ rounds and runs for time $t$; trivially, $N\leq t/\tau$. With these definitions, we can prove a lemma which will immediately yield Theorem G.2.
Lemma G.3.
Consider a single-parameter sensing problem which satisfies the HNLS criterion. Then for any initial state $\rho$,
$\|\Psi^{N}(\rho)-\Phi^{N}(\rho)\|_{1}\leq 2pN$ (175)
where $p=1-(1-\lambda)^{n}$ and $\lambda=1-e^{-4\eta\tau}$. When $\eta\tau\ll 1$, this implies
$\|\Psi^{N}(\rho)-\Phi^{N}(\rho)\|_{1}\leq 6\eta n\tau N\ .$ (176)
Proof.
Via telescoping,
$\Psi^{N}(\rho)-\Phi^{N}(\rho)=\sum_{i=0}^{N-1}\Psi^{N-i-1}\circ[(\Psi-\Phi)\circ\Phi^{i}(\rho)]\ .$ (177)
Applying the triangle inequality and using the fact that $\Psi,\Phi$ are trace-preserving,
$\|\Psi^{N}(\rho)-\Phi^{N}(\rho)\|_{1}\leq\sum_{i=0}^{N-1}\|(\Psi-\Phi)(\Phi^{i}(\rho))\|_{1}\leq N\sup_{\sigma}\|(\Psi-\Phi)(\sigma)\|_{1}$ (178)
where in the last line we absorb $\Phi^{i}$ into the state and maximize over all states. The remaining task is to bound the difference in channel output states after a single application of the channel. Note that by assumption, $\mathcal{R}$ will correct, in each round, any errors due to the signal dissipator. Only if errors due to the interface depolarizing channel occur will there be any difference between $\Psi$ and $\Phi$. We can then rewrite $\Phi=(1-p)\Psi+p\chi(\rho)$, where $\chi(\rho)$ is some quantum channel not equal to $\Psi$ and $p$ is the probability of any error due to this noise source, which we will calculate shortly. Then for any state $\sigma$,
$(\Psi-\Phi)(\sigma)=p(\Psi-\chi)(\sigma)$ (179)
and taking trace norm and using that density matrices have trace 1,
$\sup_{\sigma}\|(\Psi-\Phi)(\sigma)\|\leq p\sup_{\sigma}\|(\Psi-\chi)(\sigma)\|\leq 2p$ (180)
Precisely, $p$ is the probability that evolution by $\mathcal{L}_{dep}^{\eta}$ for time $\tau$ causes a jump of weight 1 or larger. While it is possible that the signal Lindbladian has terms which overlap with the depolarizing Lindbladian, this single-jump probability is still a valid upper bound on the channel distinguishability. Moreover, because $e^{\mathcal{L}_{\textnormal{dep}^{\eta}}\tau}=\mathcal{D}^{\otimes n}_{\lambda}$ for $\lambda=1-e^{-4\eta\tau}$, the probability of a weight $1$ or higher error is exactly $p=1-(1-\lambda)^{n}$. ∎
This trace-distance condition, as argued in [Zhou_2018], implies that the Quantum Fisher Information of the protocol increases quadratically in time only up to $\tilde{t}\propto\tau/p$, before crossing over to an asymptotic SQL sensitivity. Equivalently, attempting to resolve $\omega$ to precision $\varepsilon$ is only achievable with Heisenberg-limited sensitivity for $\varepsilon\ll 1/\tilde{t}$. For $\eta n\tau\ll 1$, these thresholds scale as $\tilde{t}=O(1/\eta n)$ and $\tilde{\varepsilon}=O(\eta n)$. This gives us Theorem G.2.
Appendix H Deferred Proofs
H.1 Proof of Lemma D.9